# Generalized Short Path Algorithms: Towards Super-Quadratic Speedup over Markov Chain Search for Combinatorial Optimization

### v1

@@ 4.2 Constructing Short Path Algorithms from Markov Chains
This subsection details how the results from [DPCB23] can be generalized to the setting of reversible Markov Chains. One of the main conditions from the aformentioned paper is that there should be a small number of low-energy states, effectively capturing that the underlying problem is hard. This is made precise through the spectral density condition, which we restate below.
Definition 4.6 ($\gamma$ Spectral Density).
$\pi(E\leq(1-\eta)E^{\ast})\leq\pi(E^{\ast})^{\gamma}.$
The tail bound given by the spectral density condition is implied by Pseudo Lipschitzness together with a functional inequality:
@@
Suppose $\mathcal{M}=(\mathcal{X},P,\pi)$ is a Markov chain with log-Sobolev constant $\omega$, and $f:\mathcal{X}\mapsto\mathbb{R}$ is $\lVert f\rVert_{P}$ pseudo-Lipschitz. Then,
@@
Suppose $\mathcal{M}=(\mathcal{X},P,\pi)$ is a Markov chain with Poincaré constant $\delta$, and $f:\mathcal{X}\mapsto\mathbb{R}$ is $\lVert f\rVert_{P}$ pseudo-Lipschitz. Then,
@@
The following corollary is immediate, and reduces spectral density to pseudo Lipschitzness and a functional inequality.
Corollary 4.9.
@@
$\mathbb{P}_{\pi}[H\leq(1-\eta)E^{\ast}]\leq\pi(E^{\ast})^{\frac{\omega((1-\eta)E^{\ast}-\mathbb{E}_{\pi}[H])^{2}}{\lVert H\rVert_{P}\ln(1/\pi(E^{\ast}))}},$
@@
Corollary 4.10.
@@
$\mathbb{P}_{\pi}[H\leq(1-\eta)E^{\ast}]\leq\pi(E^{\ast})^{\frac{\sqrt{\delta}((1-\eta)E^{\ast}-\mathbb{E}_{\pi}[H])}{\lVert H\rVert_{P}\ln(1/\pi(E^{\ast}))}},$
@@ 4.2.1 The Short Jump
The short jump is defined as the preparation of $|\psi_{b}\rangle$ from $|\sqrt{\pi}\rangle$. The ability to find a good point to short-jump to (i.e., constant $b$ where a jump takes $\operatorname{poly}(n)$ time to make) is where the inherent speedup over Grover comes from. If we just decided to do only the short-jump and sample until we found the ground state, the algorithm would only be quadratically slower due to amplitude amplification on the long-jump (assuming good gap costs $|\langle\psi_{b}|z^{*}\rangle|^{-1}$), than the full short-path algorithm.
The goal of this section is to determine conditions, using an initial Hamiltonian that is the discriminant of a reversible Markov chain, under which a short-jump can be done efficiently. However, it does not determine whether such a short-jump provides super-Grover runtime, which is the goal of the next subsection. The runtime of a short jump is captured by the short-path condition from [DPCB23, Has18a], where we present a natural generalization.
Definition 4.11 ($\theta$ Short-path condition).
Let $\Pi_{\perp}:=\mathbb{I}-|\sqrt{\pi}\rangle\!\langle\sqrt{\pi}|$. Then, the $\theta$ short-path condition holds for some constant $b>0$ if
@@
The short-path condition was used in previous papers to prove a variety of other sufficient conditions for super-Grover runtime. For example, the short path condition implies a lower bound on the spectral gap of $H_{b}$.
Lemma 4.12 ($\theta$ Short-path $\implies$ $\theta$ Spectral Gap Bound, adapted from Proposition 5 of [DPCB23]).
If $H_{b}$ satisfies the $\theta$ short-path condition, then the spectral gap of $H_{b}$ is at least $\theta$, i.e., all excited states have energy at least $-1+\theta$.
However, we stress that the short-path condition’s main purpose is to show the existence of an efficient “short jump”, i.e. $|\sqrt{\pi}\rangle\rightarrow|\psi_{b}\rangle$. In fact an inverse-$\operatorname{poly}(n)$ short-path condition is effectively equivalent to an inverse-$\operatorname{poly}(n)$ spectral gap at $H_{b}$ and overlap between $|\psi_{b}\rangle$ and $|\sqrt{\pi}\rangle$. Thus one can view it as a convenient way of combining the two conditions. The other consequence derived from the short-path condition was of a more technical nature. We summarize the implications of the short-path condition on the short jump in the following result.
@@
Suppose $\mathcal{M}=(\mathcal{X},P,\pi)$ is a reversible Markov chain with spectral gap that is at least inverse-polynomial in $n$, and $H$ is a cost Hamiltonian satisfying the $\theta$ short-path condition at $b$ independent of the problem size $n$. If $\theta=\Omega\left(\frac{1}{\operatorname{poly}(n)}\right)$, then there exists quantum algorithm for preparing an $\varepsilon$-approximation to $|\psi_{b}\rangle$ starting with $|\sqrt{\pi}\rangle$, which makes $\textup{poly}(n,\log(1/\varepsilon))$ queries to block-encodings of $D(P)$ and $H_{b}$.
@@
is inverse-polynomial in $n$. The assumption on the Markov chain gap implies that $\text{Gap}(-D(P))$ is inverse-polynomial. Lemma 4.12 implies that $\text{Gap}(H_{b})=\Omega\left(\frac{1}{\operatorname{poly}(n)}\right)$.
The overlap is implied by the following argument. Let $|\psi_{b}^{\perp}\rangle$ be the component of $\psi_{b}$ orthogonal to $|\sqrt{\pi}\rangle$. The short-path condition implies that
$\langle\psi_{b}|\Pi_{\perp}H_{b}\Pi_{\perp}|\psi_{b}\rangle\geq-1+\Omega\left(\frac{1}{\operatorname{poly}(n)}\right),$ (13)
and since $\langle\psi_{b}|H_{b}|\psi_{b}\rangle\leq-1$, we have
$\lvert\langle\psi_{b}|H_{b}|\psi_{b}\rangle-\langle\psi_{b}^{\perp}|H_{b}|\psi_{b}^{\perp}\rangle\rvert=\Omega\left(\frac{1}{\operatorname{poly}(n)}\right).$ (14)
Suppose $\lvert\langle\sqrt{\pi}|\psi_{b}\rangle\rvert=o\left(\frac{1}{\operatorname{poly}(n)}\right)$, then since $\lVert H_{b}\rVert_{2}\leq 2$
$\lvert\langle\psi_{b}|H_{b}|\psi_{b}\rangle-\langle\psi_{b}^{\perp}|H_{b}|\psi_{b}^{\perp}\rangle\rvert=o\left(\frac{1}{\operatorname{poly}(n)}\right),$ (15)
a contradiction. ∎
We will later show that an even stronger condition on the overlap can be established assuming the spectral density condition and in this case $\langle\psi_{b}|\sqrt{\pi}\rangle=1-o(1)$.
If the conditions of Theorem 4.13 are satisfied, then we only need show that $\lVert\Pi^{\ast}|\psi_{b}\rangle\rVert_{2}^{-1}$ is exponentially smaller (say by $\pi(E^{*})^{c/2}$ for some constant $c$) than $\lVert\Pi^{\ast}|\psi_{b}\rangle\rVert_{2}^{-1}$ to achieve a super-qudratic speedup over Markov chain search. This is the task of bounding the runtime of the long jump.
In this paper and [DPCB23], the long-jump runtime is upper bounded by lower bounding the easier-to-handle quantity $\lvert\langle\sqrt{\pi}|\psi_{b}\rangle\!\langle\psi_{b}|z^{*}\rangle\rvert$ by constructing an approximation to $|\psi_{b}\rangle\!\langle\psi_{b}|$. This makes use of a technical condition that ensures the ground state energy of $H_{b}$, $E_{b}$, does not decrease too much from the ground state energy of $-D(P)$. While it can be shown to hold via the short-path condition, we give intuition that it is a significantly weaker condition.
The following results provide generalized conditions under which the existence of a log-Sobolev and/or Poincaré inequality enable the spectral density condition (a very weak condition) to imply a constant $b$ at which a $\theta$ short-path condition exists, where $\theta$ is either the log-Sobolev or spectral gap of the chain $\mathcal{M}$. For Poincaré case, the spectral density condition is not even needed, although the spectral gap must of $\mathcal{M}$ must be constant.
From a bird’s-eye view, the role of the functional inequality is to upper bound a metric or divergence between $\psi_{b}^{2}$, $\ell_{2}$ distribution of $|\psi_{b}\rangle$ in the computational basis, and $\pi$. The variational definition of the corresponding metric or divergence plays the role of lower bounding. Together, these bounds are sufficient to derive a range of $b$’s where short-path holds.
@@
[…]
@@
Theorem 4.15 (Sufficient $b$ for spectral density to imply short path).
$b<\frac{2}{3}\gamma\omega\ln\left(\frac{1}{\pi(E^{\ast})}\right)$
then $\gamma$ spectral density implies an $\frac{\omega}{2}$ short-path condition.
@@
$\textup{KL}(\psi^{2}\|\pi)=\sup_{f}\{\mathbb{E}_{\psi^{2}}[f(x)]-\ln(\mathbb{E}_{\pi}[\exp(f(x))]\}.$
Choosing $f(x)=-\gamma\ln\left(\frac{1}{\pi^{\ast}}\right)g_{\eta}\left(\frac{H(x)}{E^{\ast}}\right)$ and defining $U_{\psi}=\mathbb{E}_{\psi^{2}}[g_{\eta}(\frac{H}{E^{\ast}})]$, it follows
$\textup{KL}(\psi^{2}\|\pi)$ $\geq-\gamma\ln\left(\frac{1}{\pi^{\ast}}\right)\mathbb{E}_{\psi^{2}2}\left[g_{\eta}\left(\frac{H(x)}{E^{\ast}}\right)\right]-\ln\left(\mathbb{E}_{\pi}\left[e^{-\gamma\ln\left(\frac{1}{\pi^{\ast}}\right)g_{\eta}\left(\frac{H(x)}{E^{\ast}}\right)}\right]\right)$
@@
where $\omega$ is the LS constant of $P$.
The following argument attempts to find an upper bound $b^{*}$ on the $b$, such that for all $b<b^{*}$ the two bounds above become contradicting if short-path is not satisfied.
Suppose for contradiction that the $\frac{\omega}{2}$ short-path condition is violated at $b$, i.e.
@@
Thus if short-path is violated, then the KL upper bound above reduces to
@@
We also have generally that a Poincaré inequality (LS inequality implies Poincaré) implies that for any $|\psi\rangle$
@@
where $\delta$ denotes the Poincaré constant of $P$, so if short-path is violated combining the above with Equation (22) gives
@@
where $U_{\psi_{b}^{\prime}}<0$ by construction. Also, by construction $-U_{\psi_{b}^{\prime}}\leq 1$.
@@
The proof roughly follows that of Lemma 4.14, i.e. using Equation (16), but instead uses $\pi$-Variance. From Poincaré:
@@
Note the following does not make use of the spectral density condition, which is one reason for the weak upper bound on $b$. However, it does suffice for $P$ with constant spectral gaps.
Theorem 4.17 (Sufficient $b$ for short-path under Poincaré inequality).
Suppose $\mathcal{M}=(\mathcal{X},P,\pi)$ is a reversible Markov chain that satisfies a Poincaré inequality with constant $\delta$. If
$b<\delta\frac{4\sqrt{6}-1}{10},$
then, the $\frac{\delta}{2}$ short-path condition is satisfied.
@@
$\text{TV}(\pi,\psi)=\textup{sup}_{f:\lVert f\rVert\leq 1}\frac{1}{2}(\mathbb{E}_{\psi^{2}}[f(x)]-\mathbb{E}_{\pi}[f(x)]).$
We can choose $f(x)=-g_{\eta}(H(x)/E^{*})$, which satisfies $\|f\rVert\leq 1$, so
$\text{TV}(\pi,\psi^{2})\geq\frac{1}{2}(-U_{\psi}-1),$
since $\mathbb{E}_{\pi}[f(x)]\leq 1$. Note $U_{\psi}=\mathbb{E}_{\psi^{2}}[f(x)]$.
Thus, we have the following lower bound:
$[\text{TV}(\pi,\psi^{2})]^{2}\geq\frac{1}{4}(U_{\psi}+1)^{2}.$ (27)
@@
Suppose for contradiction that the $\frac{\delta}{2}$ short-path condition is violated at $b$, i.e.
@@
where $\omega$ is the log-Sobolev constant of $P$. Thus, if $|\psi_{b}^{\prime}\rangle$ is the ground state of $\Pi_{\perp}H_{b}\Pi_{\perp}$, then
@@
Thus if short-path is violated, then the KL upper bound above reduces to
@@
If $P$ satisfies a $\delta$ Poincaré inequality, then we have that for any $|\psi\rangle$
$\langle\psi|-D(P)|\psi\rangle\geq-1+\delta,$ (31)
so if short-path is violated combining the above with Equation (29) gives
@@
where $U_{\psi_{b}^{\prime}}<0$ by construction. Also, by construction $-U_{\psi_{b}^{\prime}}\leq 1$.
@@
$0\geq\frac{1}{4}(U_{\psi_{b}^{\prime}}+1)^{2}-\frac{\delta-2bU_{\psi_{b}^{\prime}}}{2\delta}.$
As mentioned earlier, we want to a range of $b$’s that contradicts this, so we solve
$0<\frac{1}{4}(U_{\psi_{b}^{\prime}}+1)^{2}-\frac{\delta-2bU_{\psi_{b}^{\prime}}}{2\delta},$
so
$0\leq\delta U_{\psi^{\prime}_{b}}^{2}+(2\delta+4b)U_{\psi^{\prime}_{b}}-3\delta$
@@
$0\leq\delta^{2}-4\delta b-20b^{2},$
so using the positive root
$b<\delta\frac{4\sqrt{6}-1}{10}.$
Thus, the short-path condition must hold for the values of $b$ given in theorem statement. ∎
It is also natural to ask if a modified log-Sobolev inequality would work. The apparent issue is with relating the Dirichlet form to the energy with respect to $D(P)$. Although we could not prove a generic bound in terms of modified log-Sobolev constant, we can use it to derive a lower bound on standard Log Sobolev constant by using the following thoerem.
Theorem 4.18 (Theorem 1 in [Sal21]).
Let $\omega_{\textrm{LSI}}$ and $\omega_{\textrm{MLSI}}$ respectively denote the log-Sobolev constant and the modified log-Sobolev constant of a Markov chain $\mathcal{M}=(\mathcal{X},P,\pi)$. If $\mathcal{M}$ is reversible, then
$\omega_{\textrm{LSI}}\geq\omega_{\textrm{MLSI}}/\log(1/p)$

### v2

@@ 4.2 Constructing Short Path Algorithms from Markov Chains
This subsection details how the results from [DPCB23] can be generalized to the setting of reversible Markov chais. One of the main conditions from the aformentioned paper is that there should be a small number of low-energy states, effectively capturing that the underlying problem is hard. This is made precise through the spectral density condition generalized for an arbitrary chain $\mathcal{M}$ (Definition 3.2). We show that the tail bound given by the spectral density condition is implied by pseudo Lipschitzness together with a functional inequality. This follows from existing results from Markov chain theory:
@@
Suppose $\mathcal{M}=(\mathcal{X},P,\pi)$ is a Markov chain with log-Sobolev constant $\omega$, and $f:\mathcal{X}\rightarrow\mathbb{R}$ is $\lVert f\rVert_{P}$ pseudo-Lipschitz. Then,
@@
Suppose $\mathcal{M}=(\mathcal{X},P,\pi)$ is a Markov chain with Poincaré constant $\delta$, and $f:\mathcal{X}\rightarrow\mathbb{R}$ is $\lVert f\rVert_{P}$ pseudo-Lipschitz. Then,
@@
The following corollaries are immediate, and reduce spectral density to pseudo Lipschitzness and a functional inequality.
Corollary 4.8 (Log-Sobolev Inequality & Pseudo-Lipschitz $\implies\gamma$-Spectral Density).
@@
$\mathbb{P}_{\pi}[H\leq(1-\eta)E^{\ast}]\leq\pi(E^{\ast})^{\gamma},$
@@
Corollary 4.9 (Poincaré Inequality& Pseudo-Lipschitz $\implies\gamma$-Spectral Density).
@@
$\mathbb{P}_{\pi}[H\leq(1-\eta)E^{\ast}]\leq\pi(E^{\ast})^{\gamma},$
@@ 4.2.1 The Short Jump
The short jump is defined as the preparation of $|\psi_{b}\rangle$ from $|\sqrt{\pi}\rangle$. The ability to find a good point to short-jump to (i.e., constant $b$ where a jump takes $\operatorname{poly}(n)$ time to make) is where the inherent speedup over Grover comes from. If we just decided to do only the short-jump and sample until we found the ground state, the algorithm would only be quadratically slower, due to amplitude amplification on the long-jump (which assuming good gap, costs $|\langle\psi_{b}|z^{*}\rangle|^{-1}$), than the full short path algorithm.
The goal of this section is to determine conditions, using an initial Hamiltonian that is the discriminant of a reversible Markov chain, under which a short jump can be done efficiently. However, it does not determine whether such a short jump provides super-Grover runtime, which is the goal of the next subsection (Section 4.2.2). The runtime of a short jump is captured by the short path condition from [DPCB23, Has18a], where we present a natural generalization now.
Definition 4.10 ($\theta$-Short Path Condition).
Let $\Pi_{\perp}:=\mathbb{I}-|\sqrt{\pi}\rangle\!\langle\sqrt{\pi}|$. Then, the $\theta$-short path condition holds for some constant $b>0$ if
@@
The short path condition was used in previous papers to prove a variety of other sufficient conditions for super-Grover runtime. For example, the short path condition implies a lower bound on the spectral gap of $H_{b}$. {restatable}[$\theta$-Short Path Condition $\implies$ $\theta$ Spectral Gap Bound, Adapted from Proposition 5 of [DPCB23]]lemmashortPathToGapLem If $H_{b}$ satisfies the $\theta$ short path condition, then $H_{b}$ has a unique ground state and the spectral gap is at least $\theta$, i.e., all excited states have energy at least $1+\theta$.
We also show that the short-path condition implies an overlap lower bound. {restatable}[$\theta$-Short Path Condition $\implies\frac{\theta}{4}$ Overlap]lemmashortPathImpliesOverlap If $H_{b}$ satisfies the $\theta$ short path condition with $\theta\leq 1$, then $\lvert\langle\sqrt{\pi}|\psi_{b}\rangle\rvert\geq\frac{\theta}{4}$. Thus one can view it as a convenient way of combining the conditions on the gap and overlap.
We stress that the short path condition’s main purpose is to show the existence of an efficient “short jump”, i.e. transforming states $|\sqrt{\pi}\rangle\rightarrow|\psi_{b}\rangle$. The other consequence derived from the short path condition was of a more technical nature. This technical condition, called the “small-ground-state-energy shift condition”, was for bounding the cost of the long jump. Alternative conditions for this consequence of the short-path condition are presented in the next subsection.
We summarize the implications of the short path condition on the short-jump runtime in the following result.
@@
Suppose $\mathcal{M}=(\mathcal{X},P,\pi)$ is a reversible Markov chain with spectral gap that is at least inverse-polynomial in $n$, and $H$ is a cost Hamiltonian satisfying the $\theta$ short path condition at $b>0$ independent of the problem size $n$. If $\theta=\Omega\left(\frac{1}{\operatorname{poly}(n)}\right)$, then there exists quantum algorithm for preparing an $\varepsilon$-approximation to $|\psi_{b}\rangle$ starting with $|\sqrt{\pi}\rangle$, which makes $\textup{poly}(n,\log(1/\varepsilon))$ queries to block-encodings of $D(P)$ and $H_{b}$.
@@
is inverse-polynomial in $n$. The assumption on the Markov chain gap implies that $\text{Gap}(-D(P))$ is inverse-polynomial in $n$. Lemma 4.10 implies that $\text{Gap}(H_{b})=\Omega\left(\frac{1}{\operatorname{poly}(n)}\right)$, and Lemma 4.10 implies $\lvert\langle\sqrt{\pi}|\psi_{b}\rangle\rvert=\Omega\left(\frac{1}{\operatorname{poly}(n)}\right)$.
The rest of the main results of this subsection determine a range of $b$’s under which the short-path condition can be shown to hold given the $\gamma$-spectral density condition. Additionally, we provide the value of $\theta$ that is implied in each case. In Theorems 4.2 and 4.3, the $b^{\ast}$’s are a result of theorems derived in this section. Specifically, Theorems 4.14 and 4.16, provide the bounds on $b^{\ast}$ presented in Theorems 4.2 and 4.3, respectively. Theorem 4.14 shows that a log-Sobolev inequality on $\mathcal{M}$ and $\gamma$-spectral density for $H$ suffice to derive a range of $b$’s over the short-path condition is satisfied. In the case of only a Poincaré inequality and $\gamma$-spectral density, Theorem 4.16 shows that the range of such $b$’s is bounded by the Poincaré constant. In either case, $\theta$ is shown to be proportional to the LS constant or Poincaré constant, respectively. Hence, if either the LS constant or Poincaré constant is at least inverse-polynomial in $n$, then Theorem 4.14 or 4.16 plus Theorem 4.11 show that $|\psi_{b}\rangle$ can be efficiently prepared for some range of $b$’s.
It is possible that if certain conditions are not met, then $b^{\ast}$ falls with $n$, leading to an asymptotically vanishing range of $b$’s. Our results imply sufficient conditions under which we can show this does not occur. Additionally, the numerical results in Section 6 will reveal that the provable values of $b^{\ast}$ are typically overly pessimistic.
From a bird’s-eye view, the role of the functional inequality in bounding $b^{\ast}$ is to upper bound a metric or divergence between $\psi_{b}^{2}$, the $\ell_{2}$ distribution of $|\psi_{b}\rangle$ in the computational basis, and $\pi$. The variational definition of the corresponding metric or divergence is used to to establish the lower bound. Together, these bounds, via proof by contradiction, are sufficient to derive a range of $b$’s where short path condition holds.
@@
where $\textup{KL}(\cdot\|\cdot)$ denotes the Kullback–Leibler divergence, and $\psi^{2}$ is the $\ell_{2}$ distribution of $|\psi\rangle$ in the computational basis.
@@
Lemma 4.13.
Suppose the $\gamma$ spectral density condition is satisfies, then for any quantum state $\psi$
$\textup{KL}(\psi^{2}|\pi)\geq-\gamma\ln\left(\frac{1}{\pi^{\ast}}\right)\langle\psi|g_{\eta}\left(\frac{H}{E^{*}}\right)|\psi\rangle-1.$
@@
$\textup{KL}(\psi^{2}\|\pi)=\sup_{f\in\mathcal{F}{}{}}\{\mathbb{E}_{\psi^{2}}[f(x)]-\ln(\mathbb{E}_{\pi}[\exp(f(x))]\},$
where $\mathcal{F}$ denotes the set of all measurable functions. Choosing $f(x)=-\gamma\ln\left(\frac{1}{\pi^{\ast}}\right)g_{\eta}\left(\frac{H(x)}{E^{\ast}}\right)$ and defining $U_{\psi}=\mathbb{E}_{\psi^{2}}[g_{\eta}\left(\frac{H}{E^{\ast}}\right)]$, it follows
$\textup{KL}(\psi^{2}\|\pi)$ $\geq-\gamma\ln\left(\frac{1}{\pi^{\ast}}\right)\mathbb{E}_{\psi^{2}}\left[g_{\eta}\left(\frac{H(x)}{E^{\ast}}\right)\right]-\ln\left(\mathbb{E}_{\pi}\left[e^{-\gamma\ln\left(\frac{1}{\pi^{\ast}}\right)g_{\eta}\left(\frac{H(x)}{E^{\ast}}\right)}\right]\right)$
@@
Theorem 4.14 (Sufficient $b$ For Spectral Density To Imply Short Path Given Log-Sobolev).
$b<\frac{2}{3}\gamma\omega\ln\left(\frac{1}{\pi(E^{\ast})}\right)$
then $\gamma$ spectral density implies an $\frac{\omega}{2}$ short path condition.
Proof.
Let $U_{\psi}=\mathbb{E}_{\psi^{2}}[g_{\eta}\left(\frac{H}{E^{*}}\right)]$. From Lemma 4.13, we have the following lower bound on the KL divergence:
@@
where $\omega$ is the log-Sobolev constant of $P$.
The following argument attempts to find an upper bound $b^{*}$ on the $b$, such that for all $b<b^{*}$ the two bounds above become contradicting if short path is not satisfied.
Suppose for contradiction that the $\frac{\omega}{2}$ short path condition is violated at $b$, i.e.
@@
Thus if the short path condition is violated, then the $\operatorname{KL}$ upper bound above reduces to
@@
We also have generally that a Poincaré inequality (as log-Sobolev inequality implies a Poincaré inequality with the same constant) implies that for any $|\psi\rangle$
@@
where $\delta$ denotes the Poincaré constant of $P$, so if short path condition is violated combining the above with Equation (21) gives
@@
where $U_{\psi_{b}^{\prime}}<0$ by construction. Also, by construction $U_{\psi_{b}^{\prime}}\leq 1$.
@@
The proof roughly follows that of Lemma 4.12, i.e. using Equation (14), but instead uses $\pi$-Variance. From the Poincaré inequality:
@@
The following theorem provides a sufficient bound on $b$ for $\gamma$-spectral density to imply the short-path condition. There is one addtional condition relating $\pi(E^{\ast}),\gamma$ and $\delta$. However, this condition is satisfied whenever it makes sense to apply the generalized short-path algorithm. Specifically, we need the short jump to be efficient, i.e. $\delta=\Omega\left(\frac{1}{\textup{poly}(n)}\right)$ and the problem to be hard, i.e. $\pi(E^{\ast})$ is falling super polynomially.
Theorem 4.16 (Sufficient $b$ For Spectral Density To Imply Short Path Given Poincaré).
Suppose $\mathcal{M}=(\mathcal{X},P,\pi)$ is a reversible Markov chain that satisfies a $\delta<1$ Poincaré inequality. Suppose
$b<\frac{\delta}{4}.$ (26)
If the $\gamma$-spectral density is satisfied with $\frac{\delta}{2}>\pi(E^{\ast})^{\gamma}$, then the $\frac{\delta}{2}$ short-path condition is also satisfied.
@@
$\text{TV}(\pi,\psi)=\textup{sup}_{f:\lVert f\rVert_{\infty}\leq 1}\frac{1}{2}(\mathbb{E}_{\psi^{2}}[f(x)]-\mathbb{E}_{\pi}[f(x)]).$
We can choose $f(x)=-g_{\eta}(H(x)/E^{*})$, which satisfies $\|f\rVert_{\infty}\leq 1$. Let
$U_{\psi}=\mathbb{E}_{\psi^{2}}\left[g_{\eta}\left(\frac{H}{E^{*}}\right)\right].$
The $\gamma$-spectral density condition implies
$\mathbb{E}_{\pi}[f(x)]\leq\pi(E\leq(1-\eta)E^{\ast})\leq\pi(E^{\ast})^{\gamma},$
so
$\text{TV}(\pi,\psi^{2})\geq\frac{1}{2}(-U_{\psi}-\pi(E^{\ast})^{\gamma}).$ (27)
@@
Let $|\psi_{b}^{\prime}\rangle$ be the ground state of $\Pi_{\perp}H_{b}\Pi_{\perp}$ Suppose for contradiction that the $\frac{\delta}{2}$ short path condition is violated at $b$, i.e.
@@
where $\delta$ is the spectral gap of $P$. Thus
@@
Thus if short path is violated, then the total variation distance upper bound above reduces to
@@
If $P$ satisfies a $\delta$ Poincaré inequality, then we have that for any $|\psi\rangle\perp|\sqrt{\pi}\rangle$
$\langle\psi|-D(P)|\psi\rangle\geq-1+\delta.$ (31)
If short path is violated, then combining the above with Equation (29) gives
@@
where $U_{\psi_{b}^{\prime}}<0$ by construction. Also, by construction $U_{\psi_{b}^{\prime}}\leq 1$.
The above, our assumption on $\frac{\delta}{2}>\pi(E^{\ast})^{\gamma}$, assumption that $b<1$, and Equation (27) imply
$[\text{TV}(\pi,\psi^{2})]^{2}\geq\frac{1}{4}\left(-U_{\psi_{b}^{\prime}}-\pi(E^{\ast})^{\gamma}\right)^{2}.$
@@
$0\geq\frac{1}{4}(-U_{\psi_{b}^{\prime}}-\pi(E^{\ast})^{\gamma})^{2}-\frac{\delta-2bU_{\psi_{b}^{\prime}}}{2\delta}.$
As mentioned earlier, we want to find a range of $b$’s that contradicts this, so we solve
$0<\frac{1}{4}(-U_{\psi_{b}^{\prime}}-\pi(E^{\ast})^{\gamma})^{2}-\frac{\delta-2bU_{\psi_{b}^{\prime}}}{2\delta},$
which reduces to
$0<\frac{1}{4}[U_{\psi_{b}^{\prime}}^{2}+2U_{\psi_{b}^{\prime}}\pi(E^{\ast})^{\gamma}+\pi(E^{\ast})^{2\gamma}]-\frac{\delta-2bU_{\psi_{b}^{\prime}}}{2\delta}$
$0<\delta[U_{\psi_{b}^{\prime}}^{2}+2U_{\psi_{b}^{\prime}}\pi(E^{\ast})^{\gamma}+\pi(E^{\ast})^{2\gamma}]-2(\delta-2bU_{\psi_{b}^{\prime}})$
$0<\delta[-U_{\psi_{b}^{\prime}}]^{2}-(2\delta\pi(E^{\ast})^{\gamma}+4b)[-U_{\psi_{b}}^{\prime}]+\delta(\pi(E^{\ast})^{2\gamma}-2)$
@@
$0<\frac{\delta^{3}}{4b^{2}}-(2\pi(E^{\ast})^{\gamma}+4b)\frac{\delta}{2b}+\delta(\pi(E^{\ast})^{2\gamma}-2)$
$0<\delta^{3}-(2\delta\pi(E^{\ast})^{\gamma}+4b)2\delta b+\delta(\pi(E^{\ast})^{2\gamma}-2)4b^{2}$
$0<[-16+4\pi(E^{\ast})^{2\gamma}]b^{2}+4\delta\pi(E^{\ast})^{\gamma}b+\delta^{2}$
so using the positive root $b^{\ast}$:
$b^{\ast}=\frac{\delta\pi(E^{\ast})^{\gamma}}{2(4-\pi(E^{\ast})^{2\gamma})}-\frac{\sqrt{16\delta^{2}\pi(E^{\ast})^{2\gamma}-4[-16+4\pi(E^{\ast})]\delta^{2}}}{2[-16+4\pi(E^{\ast})^{2\gamma}]}$
$=\frac{\delta\pi(E^{\ast})^{\gamma}}{2(4-\pi(E^{\ast})^{2\gamma})}+4\delta\frac{\sqrt{\pi(E^{\ast})^{2\gamma}-\pi(E^{\ast})+4}}{2[16-4\pi(E^{\ast})^{2\gamma}]}$
$\geq\frac{\delta}{4}.$
It is also natural to ask if a modified log-Sobolev inequality would work. The apparent issue is with relating the Dirichlet form to the energy with respect to $D(P)$. Although we could not prove a generic bound in terms of modified log-Sobolev constant, we can use it to derive a lower bound on the standard log-Sobolev constant by using the following thoerem.
Theorem 4.17 (Theorem 1 in [STY22]).
Let $\omega_{\textup{LS}}$ and $\omega_{\textup{MLS}}$ respectively denote the log-Sobolev constant and the modified log-Sobolev constant of a Markov chain $\mathcal{M}=(\mathcal{X},P,\pi)$. If $\mathcal{M}$ is reversible, then
$\omega_{\textup{LS}}\geq\frac{\omega_{\textup{MLS}}}{\log(1/p)}$